# Fleming-Viot helps speed up variational quantum algorithms in the presence of barren plateaus

### v2

@@ Appendix B Procedure to generate 2D smooth synthetic functions
QC (1): The global minimum does not fall at a point in the border of the domain [^7] Recall footnote 6..
@@ Appendix C Method to estimate the barren percentage
Figs. 7 and 8 show the results of the barren areas classification on three function instances of each of the two optimization problem analyzed in this work; each instance corresponds to a different estimated barren level, as defined in the main text (see Section 4.1): LOW, MEDIUM and HIGH. We observe that the classification process yields reasonable results, in the sense that points classified as part of barren areas generate cohesive areas and their location tends to be consistent with contour lines.

### v3

@@ Appendix B Procedure to generate 2D smooth synthetic functions
QC (1): The global minimum does not fall at a point in the border of the domain [^6] Recall footnote 5..
@@ Appendix C Method to estimate the barren percentage
Figs. 7 and 8 show the results of the barren areas classification on three function instances of each of the two optimization problem analyzed in this work; each instance corresponds to a different estimated barren level, as defined in the main text (see Section IV-A): LOW, MEDIUM and HIGH. We observe that the classification process yields reasonable results, in the sense that points classified as part of barren areas generate cohesive areas and their location tends to be consistent with contour lines.
IEEEtran